OpenStudy (anonymous):

I will fan and medal anybody who helps me even a little bit right now, this is urgent. P(n)=-250n^2+2,500n-5,250 Part A: What are the zeroes of the above function and what do they represent? (Please show steps) Part B: Find the Maximum profit by completing the square of the function P(n). (Please show steps) Part C: What is the axis symmetry of the function P(n) Sorry if it's confusing but that's why I need help, Thank you in advance if you help! :)

OpenStudy (anonymous):

Please I really need help on this guys

OpenStudy (anonymous):

I'll Help!

OpenStudy (anonymous):

Thank you so much please do help, Im seriously so stuck on this I will medal you like a thousand times

OpenStudy (anonymous):

"A) P(n) means the profit of selling tickets at n price. So, say P(12) = 10,000. This means the profit is $10,000 if the they decide to sell the concert tickets at $12 each. find zeros means find what value of n gives P(n) = 0. suppose n is, like P(500) = 0, this means there would be no profit if each ticket cost like $500 each. Of course you to find the actual value of n that makes the equations equal 0."

OpenStudy (anonymous):

THAT should help you with the first one

OpenStudy (anonymous):

so the first section is equal to zero..?

OpenStudy (anonymous):

Oh wait okay I get it sort of

OpenStudy (anonymous):

Ok, for the second one, you just have to factor out the answer to that equation

OpenStudy (anonymous):

P(n) = -250n2 + 2,500n - 5,250 "zeros" would mean -250n^2 + 2500n - 5250 = 0 you can factor out a -250 -250 (n^2 - 10n + 21) = 0 Now factor this: n^2-10n+21, Do you know how? If not i'll help.

OpenStudy (anonymous):

Yes it would be (n-3)(n-7) right..?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now you have (n) = 3 or 7 the profitof this answer would equal to zero. This was still kinda the answer for Part A. No worries though, Part b is next :)

OpenStudy (anonymous):

And part b is much more easier :)

OpenStudy (anonymous):

Okay no problem :)

OpenStudy (anonymous):

Part B. To complete the square is use n^2 - 10n + 25 = (n-5)^2 then factor and you get n^2 - 10 + 21. so -250(n^2-10n+21+4-4) -250( (n-5)^2 - 4) -250 (n-5)^2 + 1000. so n-5=0 and solving for n you get n=5

OpenStudy (anonymous):

That's pretty much the answer for part b, and for part C, youll get your answer based on the graph to this equation and the graph looks like this...

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thank you so so much ..could you help me with ONE more with 2 parts to it? If not I will still give you a medal, you have helped me a lot :)

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Okay I will put it in a new question and tag you so you can get another medal :)

OpenStudy (anonymous):

ok :)